### v3

# MelTok:2D Tokenization for Single-Codebook Audio Compression

Abstract
Large Audio Language Models (LALMs) have emerged with strong performance across diverse audio understanding tasks and can be further enhanced by neural audio codecs. Transitioning from multi-layer residual vector quantizers to a single-layer quantizer has been shown to facilitate more efficient downstream language models decoding. However, the ability of a single codebook to capture fine-grained acoustic details remains limited, as the frequency-variant nature of 1D tokenizers leads to redundancy. To address this issue, we propose MelTok, a two-dimensional (2D) tokenizer that effectively compresses acoustic details of 44.1 KHz audio into a single codebook. The tokenizer encodes audio into a more compact representation than one-dimensional tokenizers. Furthermore, to recover audio from mel-spectrogram tokens, we propose a token-based vocoder. Both objective and subjective evaluations demonstrate that MelTok achieves quality comparable to multi-codebook codecs and outperforms existing state-of-the-art neural codecs with a single codebook on high-fidelity audio reconstruction. By preserving acoustic details, MelTok offers a strong representation for downstream understanding tasks.
Figure 1: Mel-spectrogram comparison of original and reconstructed waveforms produced by different codecs. MelTok accurately reconstructs the high-frequency details.
1 Introduction
Discrete audio tokens generated by neural audio codecs compress continuous audio signals into a compact discrete space while preserving perceptual quality and semantic content Mousavi et al. (2025). Discrete audio tokens require less storage and achieve faster transmission than continuous embeddings Theis et al. (2017). These tokens serve as an efficient and flexible interface for downstream audio understanding tasks such as Automatic Speech Recognition (ASR)  Radford et al. (2022a)  Hsu et al. (2021), sound event detection Cakir et al. (2015), music genre classification  Bahuleyan (2018), and so on. Audio codecs typically consists of an encoder-quantizer-decoder structure to encode. The encoder transforms the input waveform into a continuous representation  Langman et al. (2025), The quantizer then maps this continuous representation to a discrete code from a codebook. Finally, the decoder reconstructs the original waveform from the selected code Agustsson et al. (2017). Compression is achieved when the number of bits used to represent the code is smaller than that required for the original audio signal Yang et al. (2020).
Based on the number of quantizers used, quantization method of codecs can be broadly categorized into two types: multiple stage vector quantization  Juang and Gray (1982) and single vector quantization (SVQ). Full-Bandwidth Audio Codecs, such as DAC  Kumar et al. (2023), most commonly use residual vector quantizer (RVQ) for quantization. With iterative residual refinement, the multiple stage vector quantizer can compress 44.1 KHz audio with minimal loss in quality Kumar et al. (2023). However, multi-codebook codecs depend on multi-sequence prediction, which reduces efficiency and robustness  Li et al. (2024). Single vector quantizers has emerged as simpler and particularly useful tools for downstream generation tasks such as acoustic language models Ye et al. (2025). Recent work such as WavTokenizer  Ye et al. (2024) and UniCodec  Jiang et al. (2025) investigates compression in different domains using a single codebook. However, existing single quantizer approaches do not take “full bandwidth” rates of 44.1 or 48 kHz into account. This limitation is critical, as high-resolution acoustic detail is often required for good performance in general downstream tasks. Valin et al. (2016).
Neural audio codecs can also be categorized into two types based on the representation they compress: waveform-based and spectral-based approaches. In these waveform-based codecs, 1d tokenizers are naturally used to compress one-dimensional time-domain data Mousavi et al. (2025). DAC  Kumar et al. (2023) introduces multiscale mel reconstruction loss into this framework, which better captures details and thus improves audio quality. SNAC  Siuzdak et al. (2024) extends Residual Vector Quantization (RVQ) to multiple temporal resolutions, resulting in more efficient compression. In these waveform-based codecs, 1d tokenizers are naturally used to compress one-dimensional time-domain data.
However, high-frequency modeling is still challenging for these waveform-based models especially at high sampling rates with fewer quantizers seen in Figure  2, which is incompatible with our single-codebook objective. Spectral-based approaches solve this problem by transforming the waveform into the two-dimensional frequency domain features, which provides a more effecient representation and allows the model to better capture fine-grained frequency details Guo et al. (2025). Recent works such as Spectral CodecLangman et al. (2025) compress the mel-spectrogram and reconstruct the time-domain audio signal, which typically rely on 1D convolution kernels operating along the temporal axis. 2D convolution kernels are first used in FunCodec Du et al. (2023) after transforming waveform into spectrum, followed by a residual vector quantization (RVQ) module. While most recent codecs adopt a 1D tokenizer, they often suffer from inherent limitations: the resulting representations are not sufficiently compact. And they imperfectly model high-frequency content through one codebook, leading to audio that are clearly distinguishable from originals  Défossez et al. (2022). Our contributions are as follows:
Compression of 44.1 KHz audio in one single codebook: This codec aims to encode high-resolution audio using a single quantizer layer with minimal loss of high-frequency content, which is crucial for downstream audio understanding tasks.
Perceptual loss in mel-spectrogram reconstruction: This codec incorporates perceptual loss into mel-spectrogram reconstruction to alleviate the over-smoothing problem, and further relates it to the feature matching loss used in traditional GAN-based codecs.
A ”frequency-invariance” representation for audio: We identify a critical issue in existing single quantizer codecs which can not reconstruct full bandwidth audio due to frequency-variance redundancy. Unlike classical 1D tokenizers, which treat frequency content sequentially and are thus constrained by the assumption of frequency variance, MelTok utilizes a 2D tokenizer.
Token-based vocoder:We use a two-stage training framework to train 2D tokenizer and token-based Vocoder separately, which leads to better GAN training stability and audio quality.
Figure 2: Loss of high-frequency (above 20k Hz) detail in a waveform-based codec. Left: spectrogram of result from a waveform-based codec using 4 quantizers. Right: ground truth (GT). Noticeable differences exist in the high-frequency Mel spectra, resulting in poor reconstruction of high-frequency components, the bright ringing sound in the original sound.
2 Related Work
2.1 2D Tokenizer
Existing discrete audio tokenizers compress the waveform or spetrogram along the temporal dimension through Conv1D, so it is defined as a 1D tokenizer. A 2D tokenizer compresses 2 dimentional data in both time and frequency dimensions, and then transforms the compressed representation into a sequence of discrete tokens. Many works on 2D tokenizers have been explored in the image domain Yu et al. (2024). VQ-VAE  Yang et al. (2020) first introduced vector quantization in the latent space of VAEs to map images, audio, and video into discrete values. Vector Quantized Generative Adversarial Network (VQGAN)  Esser et al. (2021) extended this by adding perceptual and adversarial losses to better capture detail infermation. ViT-GAN further replaced convolutions with ViT Transformers  Dosovitskiy et al. (2021).
Some work in audio has also explored the usage of 2D convolutional layers. For example, PANNs  Kong et al. (2020b) employ 2D CNNs such as CNN14 to capture time–frequency invariant patterns on Wavegram features, which helps the model improves Audio Pattern Recognition performance. To explicitly capture harmonic structure, a cross-frequency structure, Ulicny et al. (2022) proposed a 2D convolution Harmonic Convolution. Inspired by the success of 2D kernels used in capturing audio structures, we aim to leverage this invariance to derive a more compact discrete representation for high-fidelity neural audio codec. Building on the structure of powerful 2D Transformer tokenizers  NVIDIA et al. (2025), we fully explore compact 2D representations for audio.
2.2 Vocoder
Neural vocoders are neural network models that converts intermediate representations, such as mel-spectrograms, into high-fidelity audio Jiao et al. (2021). Autoregressive models had long been the best-performing vocoders. WaveNet van den Oord et al. (2016), for instance, uses the mel-spectrogram as a local condition. However, its requirement for sequential (sample-by-sample) generation limits streaming efficiency. GAN-based models are capable of generating speech from mel-spectrogram efficiently Kong et al. (2020a). Since low latency is a key property for a good codec, we build our model on Vocos Siuzdak (2024). Vocos is a fast neural GAN-based vocoder designed to reconstruct audio from mel-spectrogram through inverse Fourier transform.
3 Methods
To achieve faster convergence, we propose a two-stage codec, where the first stage focuses on mel-spectrogram reconstruction with metric losses, and the second stage incorporates a discriminator to recover high-fidelity waveform from mel discrete tokens. This architecture significantly improves training efficiency, enabling our second-stage model to converge within only 50 epochs.
One efficient way to extract spectral features from an audio signal is through the Short-Time Fourier Transform (STFT). Given an input signal $x[n]$ with length T, $X_{t}[k]$, the STFT coefficient for the k-th frequency bin and the t-th time frame, denoted as $x_{t}[k]$. To better connect with the human sound perception, the frequency axis of the spectrogram can be mapped onto the Mel scale using a filter bank. This result is known as Mel spectrogram. Finally, the logarithm of the Mel spectrogram is taken to limit the range of values. The log-Mel spectrogram coefficient for the k-th frequency bin and the t-th time frame is given by:
$\underset{0\leq m\leq M-1}{LMS_{t}[m]}=\log\left[\sum_{k=0}^{N-1}H_{m}[k]\cdot\left|\sum_{n=0}^{N-1}x[n]\,w[n-tH]\,e^{-j2\pi kn/N}\right|^{2}\right].$ (1)
where $H_{m}[k]$ is the $k^{th}$ coefficient for the $m^{th}$ filter bank 200 (2001), $w[n]$ is the window function (e.g., Hamming window), H is the hop size, and N is the total number of frequency bins.
The total number of filter banks, denoted as M, determines the frequency resolution of the resulting Mel spectrogram. To preserve high-frequency details, the number of Mel filter banks M should be chosen sufficiently large, i.e., not less than 96.
3.1 2D Tokenizer VS 1D Tokenizer
Figure 3: t-SNE visualization of latent embeddings produced by the 1D and 2D tokenizers. Each point represents the latent embedding of a single time frame, and each color denotes a different frequency band.
Figure 4: Comparison of reconstructed log-mel spectrograms trained with different loss. The bottom row shows a zoomed-in view, highlighting the differences in smoothness and spectral sharpness.
Compared with conventional 1D tokenizers, our proposed 2D tokenizer provides a more compact representation for single-codebook audio compression. A 1D tokenizer processes only the temporal dimension, forcing the model to encode all frequency variation into a single token sequence. In contrast, our 2D tokenizer maintains the inherent time–frequency structure, decreasing redundancy caused by frequency variation. Furthermore, patchifying the mel-spectrogram enables the two-dimensional tokenizer to generate frequency-specific embeddings. The detail of patchify can be seen in  A.7. This leads to disentangled representations in which high-, mid-, and low-frequency structures are clearly separated in the latent space, seen in  3. These findings suggest that 2D tokenization provides a more suitable representation than 1D tokenization for single-codebook audio codecs. The encoder details are given in  7.
Figure 5: Training paradigm of MelTok.
3.2 First Stage: Mel-Spectralgram Reconstruction via 2D Tokenizer
In the first stage, we compress audio into discrete tokens and then reconstruct the mel-spectrogram from these tokens. For convenience, we adopt the log-mel representation mentioned in  equation 1, which helps constrain the value range. Our method builds on the Cosmos tokenizer  NVIDIA et al. (2025) as the foundational encoder–decoder. We optimize with the L1 loss applied on the log Mel-spectrograms, which minimizes the element-wise difference between the input and reconstructed spectrograms:
$\mathcal{L}_{\text{Mel}}=\frac{1}{T\cdot M}\sum_{t=0}^{T-1}\sum_{m=0}^{M-1}\left\lVert\mathbf{S}_{t,m}-\hat{\mathbf{S}}_{t,m}\right\rVert_{1},$ (2)
where $\mathbf{S}_{t,m}$ and $\hat{\mathbf{S}}_{t,m}$ denote the value of the input and reconstructed Mel-spectrogram, respectively, at the $t$-th time frame and the $m$-th Mel frequency bin. Using only a reconstruction loss can lead to overly smooth reconstructed Mel-spectrograms, as shown in Figure 4. This oversmoothness negatively affects downstream generation tasks such as TTS, causing the synthesized waveform to sound muffled and unnatural Sheng and Pavlovskiy (2018). In order to obtain a more detailed Mel spectrogram, we employ perceptual loss based on the VGG-19 features, given by  Simonyan and Zisserman (2015). We provide a theoretical justification for the perceptual loss on mel-spectrograms and the feature matching loss used when training generator of the vocoder in the appendix  A.5.
$\mathcal{L}_{\text{Perceptual}}=\frac{1}{L}\sum_{l=1}^{L}\sum_{t}\alpha_{l}\left\lVert\mathrm{VGG}_{l}(\hat{S})-\mathrm{VGG}_{l}(S)\right\rVert_{1},$ (3)
where $\mathrm{VGG}_{l}(\cdot)\in\mathbb{R}^{T\times M\times C}$ denotes the feature maps extracted from the $l$-th layer of a pre-trained VGG-19 network, $L$ is the number of layers considered, and $\alpha_{l}$ is the weight assigned to the $l$-th layer.
3.3 Second Stage: From Mel-Spectrogram Tokens to Waveform
Neural vocoders are primarily designed to recover audio waveforms from mel-spectrogram representations Siuzdak (2024). In contrast, our goal is to reconstruct audio waveforms from the mel-spectrogram discrete tokens obtained in the first stage, rather than from the ground-truth mel-spectrograms. Usage of these codes as input to the vocoder has advantage: it help stabilize GAN training in the second stage. Given that the first stage employs a VQ-VAE, the resulting mel-spectrograms contain reconstruction errors. As we have theoretically shown, the mapping from a continuous high-dimensional mel-spectrogram to a finite discrete codebook introduces a upper bound on the propagated error. Because the codes are discrete and belong to a finite codebook, the propagated errors are strictly bounded, preventing extreme deviations and ensuring more robust waveform recovery. By contrast, mel-spectrograms live in a continuous space, where errors cannot be strictly bounded, making the waveform recovery more sensitive to small perturbations.
3.3.1 Analysis: Bounded Error of Discrete Codes
Assumption 3.1 (Discrete Code Quantization).
Let $\mathbf{s}$ denote the original mel-spectrogram and $\mathbf{c}\in\mathcal{C}$ be the discrete token obtained from a VQ-VAE encoder $E$, where $\mathcal{C}$ is a finite codebook. We assume that the quantization error due to mapping $\mathbf{s}$ to any code $\mathbf{c}$ in codebook is bounded:
$\|\mathbf{c_{n}}-\mathbf{s}\|\leq\|\mathbf{c}-\mathbf{s}\|\leq\|\mathbf{c_{f}}-\mathbf{s}\|=\Delta,$
where $c_{f}$ denotes the farthest code, and $c_{n}$ denotes the nearest code, $\Delta$ depends on the size and codebook dimension.
Let $\mathbf{w}$ denote the waveform reconstructed in the second stage via a neural vocoder $f$. Then the reconstruction is
$\mathbf{w}=f(\mathbf{c}).$
Lemma 3.2 (Lipschitz Bound).
If $f$ is locally Lipschitz continuous with constant $L$, then
$\|f(\mathbf{c}_{1})-f(\mathbf{c}_{2})\|\leq L\|\mathbf{c}_{1}-\mathbf{c}_{2}\|,$
Theorem 3.3 (Bounded Waveform Error).
Combining Lemma 3.1 and Lemma 3.2, the error in the reconstructed waveform due to discrete code quantization is bounded:
$\|\mathbf{w}-f(\mathbf{s})\|=\|f(\mathbf{c})-f(\mathbf{s})\|\leq L\|\mathbf{c}-\mathbf{s}\|\leq L\Delta.$
Thus, the propagated error from the first-stage discrete token to the final waveform reconstruction is strictly bounded.
Assuming that the neural vocoder $f$ is $L$-Lipschitz continuous, the error propagated to the waveform $\mathbf{w}$ can be bounded by $L\Delta$. (The vocoder is a composition of convolution network, activation function and ISTFT; a proof that ISTFT is Lipschitz is given in Appendix A.6.)
3.3.2 Model Architecture
The theoretical analyse provides guidance for vocoder architecture design. For example, in the generator we choose to use the Snake activation function instead of Leaky ReLU. The Snake activation helps maintain Lipschitz continuity of the vocoder network, as its derivative is bounded by a constant of 1 Ng et al. (2025). Consequently, using Snake can help control the Lipschitz constant $L$ of the vocoder network, limiting the impact of errors from the first-stage mel-spectrogram. Also, in the discriminator we choose to use spectral normalization  Miyato et al. (2018) instead of batch normalization, which is designed to guarantee Lipschitz continuity in discriminator.
3.3.3 Training Objectives
Following the Vocos framework, our second-stage training objective consists of two key components: (i) fine-tuning the decoder from the first stage to better align the latent codes with acoustic features, and (ii) training a vocoder that translates mel-spectral codes into time-domain waveforms. To achieve this objective, we employ a combination of loss functions.
Reconstruction Loss. Reconstruction loss refers to the L1 distance between the mel-scaled magnitude spectrograms of the ground-truth waveform and the generated waveform  Kong et al. (2020a). Unlike  Yang et al. (2023) that uses 80 mel-spectrogram bins, our setup constrains the number of bins to be no smaller than the mel-spectrogram resolution defined in the first stage, which is 96 mel-spectrogram bins for music and environment sound. This ensures consistency between the first stage and second stage, preventing the loss of high-frequency details when training second stage.
Feature Matching Loss. Feature matching loss measures the learned similarity between a real and generated sample via discriminator features Larsen et al. (2016),  Kumar et al. (2019) Hifi-GAN  Kong et al. (2020a) first used it as an additional loss to train the generator of vocoder. In our case, feature matching is used to reduce over-smoothing, serving a similar role as the VGG loss applied in the first stage. Unlike the original setting in Hifi-GAN, where the feature matching loss weight is 2, we increase it to 5.
Adversarial Loss. We employ two discriminators—a multi-resolution discriminator (MRD)  Kumar et al. (2023) and a multi-period discriminator (MPD)—to enhance perceptual quality via adversarial learning Zeghidour et al. (2021).
4 Experiments
Reconstruct waveform from discrete tokens has become a fundamental task for audio codecs. In this section, we assess the performance of method relative to established baseline codecs.
DataSets. The model is trained on the AudioSet dataset, using the entire balanced training subset (bal train), musdb18-hq and the HQ-Conversations dataset Magic Data (2024). The AudioSet covers a wide range of sounds, including human and animal vocalizations, musical instruments and genres, as well as common everyday environmental noises. We evaluate the reconstruction performance of the audio domain on the AudioSet evaluation set. We evaluate the reconstruction performance of the music domain on the mixture sound in musdb18hq test set. All audio files are kept at their original sampling rate of 44 kHz. For each audio sample, mel-scaled spectrograms are computed with the following parameters: FFT size $n_{fft}$=1024, hop size $hop_{n}$ =256, and 96 Mel bins.
Training Details. In the first stage, we train the mel-spectrogram tokenizer using a combination of L1 reconstruction loss, quantization loss, and perceptual loss until convergence. Afterward, we replace the perceptual loss with a Gram-matrix loss to fine-tune the model, continuing training until convergence. During training, samples are randomly cropped to 24,320 samples, yielding a mel-spectrogram resolution of 96 × 96. We also train a vocoder using ground-truth mel-spectrograms as reference to evaluate the effect of different loss terms. In the second stage, the encoder and quantizer parameters are frozen. We jointly train the tokenizer decoder, vocoder, and discriminator.
Baseline Methods. Our proposed model is compared against DAC  Kumar et al. (2023), SNAC Siuzdak et al. (2024), Spectral Codec  Langman et al. (2025), NVIDIA NeMo Audio Codec, UniCodec  Jiang et al. (2025) and WavTokenizer  Ji et al. (2025). For all baselines, we use the officially released pretrained checkpoints— the 24kHz version for WavTokenizer and 44 kHz versions for other methods, which are publicly available online.
4.1 Evaluation
We evaluate our models using five primary objective metrics, VISQOL, LSD, Mel Distance, STFT Distance, and Mel Cepstral Distortion. We also performed subjective evaluation with MUSHRA. The primary metrics assess spectral and perceptual fidelity.
VISQOL. ViSQOL is an objective perceptual audio quality metric that compares reference and degraded audio signals to produce scores correlated with human listening judgments. In this work, we use audio mode, which operates on fullband audio at 48 kHz.
LSD. Log-Spectral Distance (LSD) is a widely used objective metric that measures the difference between the log-magnitude spectra of reference and synthesized audio, providing an indication of spectral distortion and overall reconstruction fidelity.
Mel Distance. L1 distance between the mel-scaled magnitude spectrograms of the ground truth and the generated sample.
STFT Distance. L1 distance between time-frequency representations of the ground truth and the prediction, computed using multiscale Short-Time Fourier Transform (STFT).
MCD. Mel Cepstral Distortion (MCD) is a measure of how different two sequences of mel cepstra are. The core idea is that the smaller the MCD between the original and reconstructed mel cepstral sequences, the more accurately the codec preserves the spectral characteristics and perceptual quality of the original audio signal.
For audio dataset, we use Frechet Audio Distance (FAD), which is a distribution-level metric that measures how similar reconstruted audio is to a input corpus by comparing their distributions in a deep embedding space. In our setup, we extract embeddings using a pretrained CLAP model Wu et al. (2024b) with a sample rate of 48 kHz.
For music dataset, we also use another two metrics: We calculate the similarity between the reconstruted vocal part and the original vocal part with a pre-trained wavlm-base-sv  Chen et al. (2022) and use Whisper-tiny  Radford et al. (2022a) to compute the Word Error Rate (WER) of the generated vocal part as the vocal clarity evaluation metric.
4.2 Ablation Experiment Result for Mel-spectrogram Reconstruction
To investigate the impact of 2D tokenizer used in first stage on the perceptual quality of the generated audio, we conduct an ablation study using a fixed pretrained vocoder. Specifically, we input the reconstructed mel-spectrograms obtained from 1D and 2D tokenizer into a vocoder trained with ground truth mel-spectrograms.
Table 1: Ablation Experiment One: Comparison of 1d tokenizer and 2d tokenizer used in the first stage.
[…]
1D Tokenizer 3.56 1.09 11.47 3.97 1.28
2D Tokenizer 4.36 0.67 2.64 1.68 0.48
Table 1 compares different tokenizers used in the first stage. This demonstrates that incorporating using 2d tokenizer can improves the mel-spctrogram reconstruction, which is beneficial for audio reconstruction in the second stage. We provide the encoder architecture details of 1D tokenizer and 2D tokenizer in appendix  8.
Table 2: Ablation Experiment Two: Comparison of different loss terms used in the first stage.
[…]
w/o vgg 4.18 0.76 3.82 1.87 0.57
w/ vgg 4.29 0.66 3.74 1.90 0.56
Table 2 compares different loss terms used in the first stage. This demonstrates that incorporating VGG loss effectively mitigates over-smoothing and enhances spectral reconstruction , which is beneficial for training in second stage.
4.3 Audio Reconstruction
Table 3: Objective evaluation metrics for MelTok and baselines on Audioset eval set. We bold the best results in all the models, and bold and underline the best results in single-codebook codec models. (24k: ground truth resampled to 24kHz before measurement)
[…]
DAC 9 9.0 774 0.81 2.06 0.50 4.19 3.90 0.03
SNAC 4 2.6 240 0.86 2.07 0.61 4.01 3.80 0.03
Spectral Codec 9 6.9 672 0.99 2.42 0.72 4.02 4.82 0.02
Nvidia Codec 9 6.9 672 1.07 2.45 0.74 3.96 5.39 0.08
Wavtokenizer (audio/music) 1 0.9 75 1.18 2.19 0.45(24k) 3.26 4.88 0.15
Wavtokenizer (unify) 1 0.48 40 1.24 2.37 0.52(24k) 3.26 5.32 0.16
UniCodec 1 0.9 75 1.24 2.11 0.38(24k) 3.18 4.97 0.12
Our Method 1 3.4 260 0.77 2.07 0.44 4.17 2.92 0.12
In the second stage, we use the mel-spectrogram tokens obtained from the first stage as input. After training the second stage, our final results are obtained from the jointly optimized decoder and vocoder. Our evaluation, conducted on the AudioSet test set and detailed in  3. The result shows that Mel Tok achieves competitive perceptual quality (VISQOL 4.17) while using only a single codebook. It also obtains the best fidelity metrics (lowest LSD and MCD Distance), outperforming single-codebook baseline and multi-codebook baselines.
Subject Evaluation.
We conducted a MUSHRA-style listening test to evaluate the perceptual quality of the generated audio. A total of 15 participants were recruited for the experiment. Each participant was presented with 30 randomly selected audio samples drawn from a diverse set of test cases including music, speech, and general sounds. For each trial, participants were asked to rate the audio samples on a continuous quality scale, following the MUSHRA protocol. After the test, we aggregated the ratings across all participants and samples to obtain the final statistical results. The subjective evaluation indicates that Mel Tok achieves perceptual quality comparable to other multi-codec approaches.
Figure 6: Subjective evaluation metrics calculated for different codecs. Points closer to the top-left indicate that better perceptual quality is achieved using fewer tokens, corresponding to better codec performance.
Table 4: Objective quality on the musdb18-hq test dataset.We bold the best results in all the models, and bold and underline the best results in single-codebook codec models. “Original” refers to results obtained using the ground-truth waveform. (24k: ground truth resampled to 24kHz before measurement)
[…]
DAC 0.85 1.71 0.36 4.18 2.04 1.85 0.98
SNAC 1.20 2.53 0.53 3.80 3.08 2.10 0.96
Spectral Codec 1.65 3.17 0.73 3.92 3.12 2.46 0.90
Nvidia Codec 1.77 3.24 0.76 3.84 3.28 2.22 0.92
Wavtokenizer
(audio/music)
2.01 2.20 0.47 (24k) 2.73 3.90 3.31 0.92
Wavtokenizer
2.04 2.39 0.62 (24k) 2.69 4.38 3.06 0.89
UniCodec 1.99 2.12 0.40 (24k) 2.83 3.26 2.34 0.93
Our Method 0.90 2.02 0.40 4.23 2.46 2.14 0.96
Table 5: Downstream sound event classification performance. ”Reference” refers to results obtained using the ground-truth waveform, while the other columns show the performance using audio reconstructed by different models.
Reference SNAC DAC Our Method
[…]
Table 6: Downstream music genre classification performance.
[…]
Reference 0.63 0.63 0.58
DAC 0.60 0.57 0.57
SNAC 0.63 0.59 0.57
Nvidia Codec 0.60 0.57 0.57
Spectral Codec 0.65 0.59 0.57
WavTokenizer (audio) 0.27 0.15 0.15
WavTokenizer (unify) 0.24 0.12 0.18
UniCodec 0.26 0.15 0.25
Our Method 0.57 0.50 0.50
A key property of a codec is its ability to compress and reconstruct unseen data. After augmenting the training set with hq-conversations  Magic Data (2024), we tested the model on unseen music data. The results, as reported in Table 4, demonstrate that our codec generalizes well beyond AudioSet and maintains competitive perceptual and spectral quality in out-of-distribution data.
4.4 Downstream Audio Understanding Task Evaluation
Unlike speech-only datasets, which can be evaluated using reconstructed waveform quality by ASR models, AudioSet contains multiple sound categories and is designed for audio event classification. We further evaluate our codec on this task. Specifically, we employ pretrained models from Dinkel et al. (2023) and compute the top-3 predicted labels using the reconstructed waveforms. To assess the codec’s ability to preserve semantic information, we report F1 and mAP scores in Table  5, which measure the accuracy of sound event classification. The performance demonstrates the codec’s effectiveness in retaining discriminative detail beyond perceptual quality, achieving better downstream classification results compared to other codec baselines.
Also, we performed music genre classification using a fine-tuned version of the pre-trained model DistilHuBERT  Chang et al. (2022) on the GTZAN dataset  Tzanetakis et al. (2001). As shown in Table 6, our method achieves the best performance among single-codebook codecs, while remaining competitive with multi-codebook codecs such as Spectral Codec. This demonstrates that our approach provides an effective representation for downstream music understanding tasks.
5 Limitations and Future Work
Our experiments reveal that training MelTok is disrupted when using low-resolution data that has been artificially upsampled to a high sampling rate. Currently, open-source high-fidelity audio datasets are scarce, and even 44kHz/48kHz corpora often include upsampled 16kHz content, limiting compression performance. In future work, we plan to collect more high-resolution recordings, enhance our model for better perceptual quality, and develop suitable benchmarks for evaluating non-speech, non-music sounds.
Although our experiments demonstrate that preserving more acoustic detail in the codec is promising for audio understanding tasks, its impact on downstream generation tasks remains unknown. Due to space and computational constraints, we have focused on showcasing MelTok’s reconstruction capabilities and have not yet use tokens obtained from MelTok to train in Audio Language Models (ALM). In future work, we plan to investigate the performance of MelTok-based ALMs on downstream audio tasks.
6 Conclusion
In this paper, we introduce MelTok, a neural codec that compresses full-bandwidth audio into a single codebook. We achieve this by converting mel-spectrograms into discrete tokens using a 2D tokenizer, and then reconstructing high-quality audio from these tokens with a vocoder. In the process, we investigate the disentangled distributions obtained from 2D tokenization. These findings offer new insights for future research in audio codec development. Specifically: First, incorporating perceptual loss into the reconstruction stage helps preserve more structure detail in the codebook. Second, two-stage training stabilizes the overall training process, leading to faster convergence than GAN-based models. Third, using mel-spectrograms as the representation better preserves frequency details during compression.
Appendix A Appendix
A.1 USE OF LLMs
We only used large language models (LLMs) as a tool for language refinement and editing. They were not involved in the design of the methodology, experimental setup, data analysis, or any other core aspect of this research.
A.2 ETHICS STATEMENT
This study has been approved by the relevant ethics committee or institutional review board and was conducted in strict accordance with ethical guidelines. The rights, privacy, and welfare of participants were fully respected and protected, and all personal information was kept confidential.
Informed Consent: All participants were informed of the study’s objectives, procedures, potential risks, and benefits, either verbally or in writing, and provided their informed consent.
Data Confidentiality and Privacy Protection: Measures were implemented to safeguard participants’ personal information and ensure privacy.
A.3 REPRODUCIBILITY STATEMENT
Use of Research Data: All research data were collected, stored, and used in accordance with legal and ethical standards, ensuring transparency and proper interpretation.
We ensure that our method is fully reproducible. Upon acceptance of this paper, we will publicly release all code, model weights, and training data necessary to replicate our results.
A.4 Equivalence of Spectrogram VGG Loss and Vocoder Feature Matching Loss
Let $\hat{x}$ denote the waveform generated by a vocoder, and $x$ the corresponding ground-truth waveform. Define the Short-Time Fourier Transform (STFT) of the waveform as:
$\hat{S}=\mathrm{STFT}(\hat{x}),\quad S=\mathrm{STFT}(x)$ (4)
A.4.1 Spectrogram VGG Loss
A spectrogram-based VGG loss is defined as the L1 distance between feature maps extracted from a convolutional network $\phi$ (e.g., VGG) applied to the spectrograms:
$\mathcal{L}_{\mathrm{VGG}}(\hat{S},S)=\sum_{l=1}^{L}w_{l}\,\|\phi_{l}(\hat{S})-\phi_{l}(S)\|_{1}$ (5)
where $\phi_{l}(\cdot)$ is the feature map at the $l$-th layer, $w_{l}$ is a weighting coefficient, and $L$ is the total number of layers considered.
A.4.2 Vocoder Feature Matching Loss
In vocoder GANs, the feature matching loss is defined using the discriminator $D$:
$\mathcal{L}_{\mathrm{FM}}(\hat{x},x)=\sum_{d\in\mathcal{M}}\sum_{l=1}^{L_{d}}\|D_{l}^{(d)}(\hat{x})-D_{l}^{(d)}(x)\|_{1}$ (6)
where $D_{l}^{(d)}(\cdot)$ denotes the feature map of the $l$-th layer of the $d$-th discriminator, and $\mathcal{D}$ is the set of discriminators (e.g., multi-resolution discriminators).
Each discriminator first computes a spectrogram of the waveform:
$X=\mathrm{STFT}(\cdot)$ (7)
and then applies a sequence of convolutional layers with non-linearities:
$D_{l}^{(d)}(\hat{x})=\sigma(W_{l}^{(d)}*X+b_{l}^{(d)}),$ (8)
where $W_{l}^{(d)},b_{l}^{(d)}$ are the convolutional weights and biases, and $\sigma(\cdot)$ is the activation function.
A.4.3 Equivalence
[…]
$\mathcal{L}_{\mathrm{FM}}(\hat{x},x)=\sum_{d,l}\Big\|\sigma(W_{l}^{(d)}*\mathrm{STFT}(\hat{x})+b_{l}^{(d)})-\sigma(W_{l}^{(d)}*\mathrm{STFT}(x)+b_{l}^{(d)})\Big\|_{1}$ (9)
Comparing with the spectrogram VGG loss in Eq. (2), we see that the two losses share the same mathematical form:
$\sum_{l}\|F_{l}(\mathrm{STFT}(\hat{x}))-F_{l}(\mathrm{STFT}(x))\|_{1}$ (10)
where $F_{l}$ denotes a convolutional feature extractor. The only difference lies in the choice of network parameters (pretrained VGG weights vs. learned discriminator weights).
A.5 Equivalence of VGG Loss and Feature-Matching Loss
Both the spectrogram-based VGG loss and the vocoder feature matching loss are equivalent in the sense that they compute an L1 distance in the convolutional feature space of a spectrogram. Formally,
$\mathcal{L}_{\mathrm{VGG}}(\hat{S},S)\approx\mathcal{L}_{\mathrm{FM}}(\hat{x},x),$ (11)
up to the network weights. This shows that the vocoder feature matching loss can be interpreted as a generalized, learnable spectrogram-based perceptual loss.
A.6 Proof of ISTFT Lipschitz Continuity
We prove that the inverse short-time Fourier transform (ISTFT) operator is Lipschitz continuous with respect to the spectrogram input.
Lemma A.1 (ISTFT Lipschitz Continuity).
Let $\mathcal{S}\in\mathbb{C}^{F\times T}$ be a complex spectrogram obtained by short-time Fourier transform (STFT) with analysis window $g\in\mathbb{R}^{N}$ and hop size $H$. Define the ISTFT operator $\mathrm{ISTFT}:\mathbb{C}^{F\times T}\to\mathbb{R}^{M}$ with synthesis window $h$. Then, for any two spectrograms $\mathcal{S}_{1},\mathcal{S}_{2}$,
$\|\mathrm{ISTFT}(\mathcal{S}_{1})-\mathrm{ISTFT}(\mathcal{S}_{2})\|_{2}\;\leq\;L_{\mathrm{ISTFT}}\,\|\mathcal{S}_{1}-\mathcal{S}_{2}\|_{2},$
where the Lipschitz constant $L_{\mathrm{ISTFT}}$ depends only on the window functions and hop size.
Proof.
Recall that ISTFT reconstructs the waveform by overlap-add (OLA) of inverse FFTs of each frame:
$\hat{x}[n]=\sum_{t}h[n-tH]\cdot\mathrm{IFFT}(\mathcal{S}[:,t])[n-tH].$
Let $\Delta\mathcal{S}=\mathcal{S}_{1}-\mathcal{S}_{2}$. Then the waveform difference is
$\Delta x[n]=\sum_{t}h[n-tH]\cdot\mathrm{IFFT}(\Delta\mathcal{S}[:,t])[n-tH].$
By Parseval’s theorem, the $\ell_{2}$ norm of the IFFT is equal to $\ell_{2}$ norm of the spectrum:
$\|\mathrm{IFFT}(\Delta\mathcal{S}[:,t])\|_{2}=\sqrt{N}\,\|\Delta\mathcal{S}[:,t]\|_{2},$
where $N$ is the FFT length.
Applying Minkowski’s inequality to the OLA sum:
$\|\Delta x\|_{2}\leq\sum_{t}\|h(\cdot-tH)\|_{\infty}\cdot\|\mathrm{IFFT}(\Delta\mathcal{S}[:,t])\|_{2}.$
Since the shifted window has the same maximum magnitude as $h$,
$\|\Delta x\|_{2}\leq\|h\|_{\infty}\cdot\sqrt{N}\,\sum_{t}\|\Delta\mathcal{S}[:,t]\|_{2}.$
Finally, by Cauchy–Schwarz,
$\sum_{t}\|\Delta\mathcal{S}[:,t]\|_{2}\leq\sqrt{T}\,\|\Delta\mathcal{S}\|_{2}.$
Combining the inequalities, we obtain
$\|\mathrm{ISTFT}(\mathcal{S}_{1})-\mathrm{ISTFT}(\mathcal{S}_{2})\|_{2}\;\leq\;\|h\|_{\infty}\cdot\sqrt{NT}\,\|\mathcal{S}_{1}-\mathcal{S}_{2}\|_{2}.$
Thus ISTFT is Lipschitz continuous with constant
$L_{\mathrm{ISTFT}}\leq\|h\|_{\infty}\cdot\sqrt{NT}.$
Remark A.2.
In practice, when $h$ is chosen as the canonical synthesis window satisfying the perfect-reconstruction condition (e.g., Hann window with $50\%$ overlap), $\|h\|_{\infty}\leq 1$. Hence the ISTFT operator has a moderate Lipschitz constant that scales with the FFT length $N$ and number of frames $T$, ensuring stability against spectrogram perturbations.
A.7 Architecture Details For Tokenization
Table 7: The encoder architecture details of MelTok
Layer Name
Input Shape
(Batch,Channel,Frequency,Time)
Kernel Size, Stride Output Shape
Patcher (8, 1, 96, 96) (8, 1, 48, 48)
PreConv2D (8, 1, 48, 48) (3, 1) (8, 128, 48, 48)
Downsampling 1:
ResnetBlock (8, 128, 48, 48) (3, 1), (1, 1) (8, 256, 48, 48)
ResnetBlock (8, 256, 48, 48) (3, 1), (1, 1) (8, 256, 48, 48)
DownConv2D (8, 256, 48, 48) (2, 2) (8, 256, 24, 24)
AttnBlock (Cond.) (8, 256, 24, 24) (1, 1) (8, 256, 24, 24)
Downsampling 2:
ResnetBlock (8, 256, 24, 24) (3, 1), (1, 1) (8, 512, 24, 24)
ResnetBlock (8, 512, 24, 24) (3, 1), (1, 1) (8, 512, 24, 24)
DownConv2D (8, 512, 24, 24) (2, 2) (8, 512, 12, 12)
AttnBlock (Cond.) (8, 512, 12, 12) (1, 1) (8, 512, 12, 12)
ResnetBlock $\times$ 2 (8, 512, 12, 12) (3, 1), (1, 1) (8, 512, 12, 12)
AttnBlock (Cond.) (8, 512, 12, 12) (1, 1) (8, 512, 12, 12)
Normalizer (8, 512, 12, 12) (8, 512, 12, 12)
PostConv2D (8, 512, 12, 12) (3, 1) (8, 512, 12, 12)
Table 8: 2D Tokenizer vs 1D Tokenizer Architecture in Ablation Study
2D Tokenizer 1D Tokenizer
Input $\times 96 96×96$$\times 96$
Encoder $96\times 96\to 4\times 48\times 48$ (Haar, patch=2) $96\times 96\to 128\times 96$
$4\times 48\times 48\to 128\times 48\times 48$ $128\times 96$$256\times 96$
Downsample: Downsample:
$128\times 48\times 48\to 256\times 24\times 24\to 512\times 12\times 12$ $256\times 96\to 512\times 12$
Pre quant conv $512\times 12\times 12\to 256\times 12\times 12$ $512\times 12\to 256\times 12$
Quant_conv $256\times 12\times 12\to 64\times 12\times 12$ $256\times 12\to 64\times 12$
Quantizer 8192 codes of dim 64 8192 codes of dim 64
One Quantizer 12 Quantizers to keep same bitate
Post quant conv $64\times 12\times 12\to 256\times 12\times 12$ $64\times 12\to 256\times 12$
Decoder Symmetric to Encoder Symmetric to Encoder
Reconstruction $\times 96 96×96$$\times 96$